### v5

# Tiny-BioMoE: a Lightweight Embedding Model for Biosignal Analysis

Abstract.
Pain is a complex and pervasive condition that affects a significant portion of the population. Accurate and consistent assessment is essential for individuals suffering from pain, as well as for developing effective management strategies in a healthcare system. Automatic pain assessment systems enable continuous monitoring, support clinical decision-making, and help minimize patient distress while mitigating the risk of functional deterioration. Leveraging physiological signals offers objective and precise insights into a person’s state, and their integration in a multimodal framework can further enhance system performance. This study has been submitted to the Second Multimodal Sensing Grand Challenge for Next-Gen Pain Assessment (AI4PAIN). The proposed approach introduces Tiny-BioMoE, a lightweight pretrained embedding model for biosignal analysis. Trained on $4.4$ million biosignal image representations and consisting of only $7.3$ million parameters, it serves as an effective tool for extracting high-quality embeddings for downstream tasks. Extensive experiments involving electrodermal activity, blood volume pulse, respiratory signals, peripheral oxygen saturation, and their combinations highlight the model’s effectiveness across diverse modalities in automatic pain recognition tasks. The model’s architecture (code) and weights are available at https://github.com/GkikasStefanos/Tiny-BioMoE.
Pain assessment, pain recognition, deep learning, multimodal, foundation model, data fusion
journalyear: 2025doi: XXXXXXX.XXXXXXXconference: October 13–17, 2025; Canberra, Australia; $c$s: Applied computing Health informatics
1. Introduction
Pain serves as a vital evolutionary mechanism, alerting the organism to potential harm or signaling the onset of illness, and plays a vital role in the body’s defense system by helping maintain physiological integrity (Santiago, 2022). It is a subjective experience comprising multiple components, including nociceptive, sensory, affective, and cognitive dimensions (Marchand, 2024). Pain has been described as a “Silent Public Health Epidemic” (Katzman and Gallagher, 2024), emphasizing its widespread yet often underestimated impact. In nursing literature, it is also referred to as “the fifth vital sign” (Joel, 1999), reflecting the need for its routine and systematic assessment alongside other vital signs. In addition, opioid analgesics are the most frequently prescribed treatment for pain management (Kaye et al., 2017), yet they often lead to addiction and overdose (Stampas et al., 2020). Moreover, their side effects—such as lethargy, depression, anxiety, and nausea—significantly affect both workforce productivity and overall quality of life (Benyamin et al., 2008). The inherent subjectivity and complexity of pain assessment have been identified as significant challenges in both research and clinical practice. For example, pain management often relies on patients’ subjective reports, making it difficult to administer medication with precision. This lack of objective evaluation contributes significantly to the overprescription and overuse of pain medications (Kong and Chon, 2024). Moreover, managing and assessing pain in patients with–or at risk of–medical instability presents significant clinical challenges, particularly when communication barriers are present (Puntillo et al., 2002). Research highlights that pain in critically ill adults remains frequently under-managed. A significant limitation is the lack of structured, comprehensive tools for assessing pain and supporting clinical decision-making in these contexts (Meehan et al., 1995). Furthermore, cancer-related pain is highly prevalent, particularly in advanced stages of the disease, with its incidence exceeding $40\%$ (Bang et al., 2023). Additionally, the presence of negative emotions such as fear and sadness can influence pain expression, further complicating accurate assessment (Tessier et al., 2024).
Precise evaluation and understanding of the factors influencing pain are crucial for achieving effective pain management (Sabbadini et al., 2020). Self-report methods, such as numerical rating scales and questionnaires, remain the gold standard for evaluating patient experiences. However, their reliability significantly diminishes in cases where patients exhibit altered consciousness, cognitive impairments, or communication difficulties (Herr et al., 2006). Behavioral cues—such as facial expressions, vocalizations, and body movements—are widely used to infer pain, particularly in patients who cannot communicate (Fernandez Rojas et al., 2023a). Complementary to these, physiological signals like electrocardiography (ECG), electromyography (EMG), and electrodermal activity (EDA) provide deeper insights into the body’s response to pain (Gkikas and Tsiknakis, 2023a). Physiological signals play a crucial role in pain assessment by enabling a more objective and accurate understanding of a person’s condition. Sabbadini et al. (Sabbadini et al., 2024) emphasized the importance of incorporating raw physiological data, such as continuous waveforms, to capture the intricate relationship between pain and physiological responses. Gozzi et al. (Gozzi et al., 2024) advocated for a broader perspective on pain as a multidimensional phenomenon, emphasizing the integration of physiological biomarkers with psychosocial factors.
This study introduces a lightweight embedding model, pretrained on over $4$ million biosignals, and applies it to automatic pain assessment tasks across a wide range of modalities. Although numerous studies rely on biosignals, few models leverage large-scale pretraining. Moreover, given the growing concerns around the computational cost of large models for both training and inference, this work proposes a compact and efficient alternative, aiming to ensure accessibility regardless of the user’s hardware capabilities.
2. Related Work
Over the past decade, automatic pain assessment has gained increasing attention, with advancements shifting from conventional image and signal processing methods to more complex deep learning-based techniques (Gkikas, 2025). The majority of existing methods are video-based, aiming to capture behavioral cues through facial expressions, body movements, or other visual indicators and employing a wide range of modeling strategies (Gkikas and Tsiknakis, 2023b; Bargshady et al., 2024; Gkikas and Tsiknakis, 2024a; Huang et al., 2022). While video-based approaches dominate the field, a considerable number of studies have also focused on biosignal-based methods, although to a lesser extent. These works have investigated the utility of various physiological signals, such as electrocardiography (ECG) (Gkikas. et al., 2022; Gkikas et al., 2023), electromyography (EMG) (Pavlidou and Tsiknakis, 2025; Patil and Patil, 2024; Thiam et al., 2019; Werner et al., 2014), electrodermal activity (EDA) (Aziz et al., 2025; Li et al., 2025; Lu et al., 2023; Phan et al., 2023; Ji et al., 2023), and brain activity through functional near-infrared spectroscopy (fNIRS) (Rojas et al., 2016; Fernandez Rojas et al., 2019; Rojas et al., 2021; Fernandez Rojas et al., 2024b; Khan et al., 2024; Bargshady et al., 2025). For a more comprehensive analysis of biosignal modalities within automatic pain recognition frameworks, the reader is referred to (Khan et al., 2025b).
In addition, multimodal approaches combining behavioral and physiological data have gained increasing attention in recent years, with several studies demonstrating the benefits of integrating multiple sources of information to improve performance. Many researchers explore combination of biosignals due to their potential to provide more accurate measurements. For instance, Jiang et al. (Jiang et al., 2024a) proposed a hybrid temporal-channel attention model that fuses ECG and galvanic skin response (GSR) signals, while in (Jiang et al., 2024b), the authors employed ECG and GSR signals alongside neural networks enhanced with Squeeze-and-Excitation blocks. Chu et al. (Chu et al., 2017) integrated blood volume pulse, ECG, and skin conductance level, using a hybrid approach that combined genetic algorithm-based feature selection with principal component analysis. On the other hand, Badura et al. (Badura et al., 2021) employed a multimodal setup in their effort to assess pain during physiotherapy, which included EDA, EMG, respiration, blood volume pulse (BVP), and hand force. Other researchers focus on combining physiological with behavioral modalities. The authors in (Zhi and Yu, 2019) utilized facial videos in combination with ECG, EMG, and EDA, extracting a wide range of handcrafted features and fusing them. In (Gkikas et al., 2024), the authors combined facial videos and heart rate data to develop a transformer-based framework for recognizing pain intensity. Similarly, in (Farmani et al., 2025), CNN and LSTM models were used to analyze facial videos and EDA, achieving high performance. Finally, the authors in (Gkikas et al., 2025c) proposed a foundation model that extracted feature representations from a wide range of modalities—both behavioral and physiological—for pain assessment.
3. Methodology
This section outlines the architecture of the proposed embedding model, the pretraining procedure, and the dataset used for pretraining. It also details the biosignal preprocessing steps and the transformation of these signals into visual representations. Additionally, it describes the augmentation and regularization techniques employed for the pain recognition tasks.
3.1. Tiny-BioMoE
Inspired by recent advancements in deep learning involving Mixture of Experts (MoE) architectures for language (Cai et al., 2025) and vision tasks (Riquelme et al., 2021), we propose Tiny-BioMoE—a lightweight MoE model comprising only $7.34$ million parameters. The model consists of two vision transformer encoders, Encoder-1 and Encoder-2. Both encoders process the input image independently to extract their respective embedding representations, which are subsequently fused into a unified feature vector. Table 2 reports the number of parameters and computational cost, measured in floating-point operations (FLOPs), while Figure 1 provides a high-level architectural overview.
Figure 1. Overview of the Tiny-BioMoE architecture.
3.1.1. Encoder-1
Integrates spectral and self-attention mechanisms in the initial layers, while relying exclusively on self-attention in the later stages. Each input image $I$ is divided into $n$ non-overlapping $16\times 16$ patches, reshaped into tokens $\in\mathbb{R}^{n\times d}$ with $d=768$, followed by positional encoding. To incorporate frequency information, a 2D Fast Fourier Transform (FFT) is applied to each token $x$, yielding $X=\mathscr{F}[x]\in\mathbb{C}^{h\times w\times d}$. A learnable complex-valued filter $K\in\mathbb{C}^{h\times w\times d}$ modulates the spectral components via element-wise multiplication: $\tilde{X}=K\odot X$. The filtered signal is then transformed back into the spatial domain using the inverse FFT: $x\leftarrow\mathscr{F}^{-1}[\tilde{X}$]. A depthwise convolution-based MLP enhances channel-wise interactions, $f(x)=W_{2}\cdot\text{GELU}(\text{DWConv}(W_{1}\cdot x+b_{1}))+b_{2}$, with layer normalization applied before and after the FFT and IFFT. The attention layer employs standard scaled dot-product self-attention:
(1) $\widetilde{X}=\text{Attn}(XW_{q},XW_{k},XW_{v}),$
where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d}$ are the projection matrices for queries, keys, and values, respectively. Encoder-1 is composed of four hierarchical stages, each reducing spatial resolution by a factor of 2. The embedding dimensions across the stages are $64$, $128$, $320$, and $96$, with $1$ attention head used per stage. Figure 2(a) illustrates the architecture of Encoder-1.
3.1.2. Encoder-2
Built upon hierarchical vision-transformer principles, incorporates mechanisms that enhance both efficiency and speed. Encoder-2 consists of two core components: the Spatial-Mixer and the Waterfall-Attention module. The architecture is organized with Waterfall-Attention at its center, preceded and followed by Spatial-Mixer modules. The input image $I$ is first divided into overlapping $16\times 16$ patches using a projection layer, with each patch embedded into a $d$-dimensional token. To capture local information, each patch $T$ is processed with depth-wise convolution:
(2) $Y_{c}=K_{c}*T_{c}+b_{c},$
where $T_{c}$ and $Y_{c}$ are the input and output of channel $c$, $K_{c}$ is the channel-specific kernel, and $b_{c}$ is the corresponding bias. Following convolution, batch normalization is applied as $Z_{c}=\text{BN}(Y_{c})$, where BN denotes channel-wise normalization with learnable affine parameters. A feed-forward network (FFN) then facilitates inter-channel communication, $\Phi^{F}(Z_{c})=W_{2}\cdot\text{ReLU}(W_{1}\cdot Z_{c}+b_{1})+b_{2}$, where $\Phi^{F}(Z_{c})$ is the output of the feed-forward network for the input $Z_{c}$. $W_{1}$ and $W_{2}$ are the weight matrices of the first and second linear layers; $b_{1}$ and $b_{2}$ are the bias terms for the first and second linear layers, respectively, and ReLU is the activation function. Encoder-2 uses a single self-attention layer. For each input embedding, $X_{i+1}=\Phi^{A}(X_{i})$ is computed, where $X_{i}$ is the full input embedding for the $i$-th block. The Waterfall-Attention module partitions the embedding into $h$ segments, assigns each to a distinct head, and distributes the workload:
(3) $\widetilde{X}_{ij}=Attn({X}_{ij}W^{Q}_{ij},{X}_{ij}W^{K}_{ij},{X}_{ij}W^{V}_{ij}),$
(4) $\widetilde{X}_{i+1}=\text{Concat}[\widetilde{X}_{ij}]_{j=1:h}W^{P}_{i},$
where each $j$-th head processes $X_{ij}$, the $j$-th segment of the full embedding $X_{i}$, structured as $X_{i1},X_{i2},\dots,X_{ih}]$. Projection layers $W^{Q}_{ij}$, $W^{K}_{ij}$, and $W^{V}_{ij}$ map each segment into distinct subspaces, and $W^{P}_{i}$ reunites the concatenated outputs. The waterfall design enriches each subsequent head by residual addition: $X^{{}^{\prime}}_{ij}=X_{ij}+\widetilde{X}_{i(j-1)}$. Depth-wise convolution on every $Q$ enables the self-attention to fuse global context with local cues. The model stacks three stages, each one a token layer deep. It halves the spatial resolution at every stage, progressively reducing the number of tokens. The associated embedding widths are $d=192$, $288$, and $96$, and the stages employ $3$, 33, and $4$ attention heads, respectively. Figure 2(b) illustrates the architecture of Encoder-2.
Figure 2. Representation of the architecture of the main components of Tiny-BioMoE: (a) Encoder-1 and (b) Encoder-2.
3.1.3. Fusion
Initially, a LayerNorm operation is applied to the input image tensor to ensure consistent normalisation before processing. The normalised input is then passed to both Encoder-1 and Encoder-2, producing embeddings $z_{1}$ and $z_{2}$, respectively. Each output is further normalised using a second LayerNorm, yielding $\hat{z}_{1}$ and $\hat{z}_{2}$. A lightweight gating network is defined as:
(5) $g(x)=\text{HardTanh}(\text{ELU}(Wx)),$
and produces per-channel modulation coefficients $\alpha_{1},\alpha_{2}\in[0,1]$. These are applied element-wise to the encoder outputs:
(6) $z^{\prime}_{1}=\alpha_{1}\odot\hat{z}_{1},\quad z^{\prime}_{2}=\alpha_{2}\odot\hat{z}_{2}.$
Each encoder produces a $96$-dimensional embedding, and after re-weighting and concatenation, the resulting feature vector becomes $192$-dimensional:
(7) $z_{\text{cat}}=[z^{\prime}_{1}\,\|\,z^{\prime}_{2}]\in\mathbb{R}^{192}.$
A final LayerNorm is applied to $z_{\text{cat}}$, yielding the unified output used in downstream tasks.
3.2. Biosignal Pre-processing & Visualization
The biosignal samples used for both pretraining and pain-related tasks were converted into visual representations to enable compatibility with vision-based models. In total, six distinct transformations were applied: (1) Spectrogram-Angle plots encode the phase angle of the frequency components; (2) Spectrogram-Phase includes phase information with unwrapping applied to resolve discontinuities; (3) Spectrogram-PSD displays the power spectral density, reflecting how signal power varies across frequencies and time; (4) Recurrence plots map the recurrence of states in the signal’s phase space to reveal temporal dynamics; (5) Scalograms represent time-frequency content using continuous wavelet transforms; (6) Waveform diagrams visualize the raw signal over time, capturing its amplitude, frequency, and phase characteristics. Figure 3 shows one example of each of these six visual representations. In addition, only the biosignals used in the pain-related tasks i.e., electrodermal activity, blood volume pulse, respiratory signals, and peripheral oxygen saturation were filtered using a low-pass $5$ Hz and band-pass filters of 0.041.70.04-1.7 Hz, 0.050.50.05-0.5 Hz, and 0.041.70.04-1.7 Hz, respectively.
Figure 3. Examples of the six visual representations used for biosignals. From left to right: Spectrogram-Angle, Spectrogram-Phase, Spectrogram-PSD, Recurrence plot, Scalogram, and Waveform.
3.3. Pretraining
Tiny-BioMoE, the proposed model, serves as an embedding extractor for biosignals. It was trained on $3$ large-scale datasets comprising a total of $4.4$ million samples—details are provided in Table 1. These datasets cover a wide range of biosignal modalities, including EEG, EMG, and ECG. The model was trained using a multi-task learning framework, where each dataset is considered as a distinct supervised task. All $14$ tasks are learned jointly using the following objective:
(8) $L_{\text{total}}=\sum_{i=1}^{3}\left[e^{w_{i}}L_{S_{i}}+w_{i}\right],$
where $L_{S_{i}}$ is the loss associated with the $i$-th task, and $w_{i}$ are trainable weights that adaptively control each task’s contribution to the overall objective. Training was conducted for $200$ epochs under this setup.
Table 1. Datasets utilized for the multitask learning-based pretraining process of the PainFormer.
Dataset Number of Samples Modality
EEG-BST-SZ (Ford et al., 2013) 1.20M EEG
Silent-EMG (Gaddy and Klein, 2020) 1.03M EMG
ECG HBC Dataset (Kachuee et al., 2018) 2.16M ECG
Total: 3 datasets–tasks 4.39M
  • The reported number of samples refers exclusively to the training split, excluding validation and testing sets. Each sample in the corresponding dataset is represented by six distinct visualizations, as described in 3.2.
Table 2. Number of parameters and FLOPS for the components of the proposed Tiny-BioMoE.
Module Parameters (M) FLOPS (G)
Encoder-1 2.90 2.36
Encoder-2 4.13 0.68
Tiny-BioMoE 7.34 3.04
3.4. Augmentation Methods & Regularization
Several data augmentation techniques are applied during training for the pain recognition tasks. Every $224\times 224$ image undergoes a cascade of stochastic transformations. The AugMix method blends three randomly generated augmentation chains with the original image, shifting a mixture of contrast, colour, and geometric perturbations to the image. In addition, TrivialAugment applies a single randomly chosen operation with a magnitude sampled uniformly. Centre cropping is applied with a probability drawn from a given range, where the crop size is selected randomly and the image is resized back to its original dimensions. Conditional Gaussian blurring applies noise by reducing high-frequency components. Two Cutout masks are applied—one placing a small number of blocks, the other covering the image with a higher number, both using blocks of equal size, $32\times 32$. Beyond data-level augmentation, two stochastic regularizers have been utilized. A Dropout layer follows the encoder, with its keep probability gradually reduced linearly over training epochs:
(9) $p(t)\;=\;p_{\text{start}}+\frac{t}{T}\bigl{(}p_{\text{end}}-p_{\text{start}}\bigr{)},\qquad 0\leq t\leq T.$
The cross-entropy loss employs a Label-Smoothing based on the same linear schedule, shifting from a soft target distribution toward one-hot labels. Finally, a cosine learning-rate profile with Warmup and Cooldown schedules is also employed. All stochastic choices—operation selection, magnitudes, cropping ratios, mask locations, and probability draws—are resampled independently for every image at every batch. Throughout all experiments, the batch size is fixed to $32$ and the learning rate is set to $1\mathrm{e}{-4}$. Figure 4 presents an overview of the proposed multimodal pipeline.
Figure 4. Schematic overview of the proposed pipeline for pain assessment using various modalities and visual representation of them.
4. Experimental Evaluation & Results
This study leverages the dataset released by the challenge organizers, which consists of electrodermal activity recordings from $65$ participants. Data collection took place at the Human-Machine Interface Laboratory, University of Canberra, Australia, and is divided into $41$ training, $12$ validation, and $12$ testing subjects. Pain stimulation was induced using transcutaneous electrical nerve stimulation (TENS) electrodes positioned on the inner forearm and the back of the right hand. Two pain levels were measured: pain threshold—the minimum stimulus intensity perceived as painful (low pain), and pain tolerance—the maximum intensity tolerated before becoming unbearable (high pain). The signals have a frequency of $100$ Hz and a duration of approximately $10$ seconds. We refer to (Fernandez Rojas et al., 2023b) for a detailed description of the recording protocol and to (Fernandez Rojas et al., 2024a) for information regarding the previous edition of the challenge. This study utilises all available modalities—EDA, BVP, respiratory signals, and peripheral oxygen saturation (SpO2). All experiments were conducted on the validation subset of the dataset and evaluated within a multi-class classification framework, encompassing three pain levels: No Pain, Low Pain, and High Pain. The validation results are reported in terms of macro-averaged accuracy, precision, and F1 score. The final results of the testing set are also reported. We note that all experiments followed a deterministic setup, eliminating the effect of random initializations; thus, any performance differences arose strictly from the chosen optimization settings, modalities, or other intentional changes rather than chance.
4.1. The Impact of Pretraining
The first series of experiments evaluates the impact of pretraining on model performance. All experiments followed a consistent training setup of $200$ epochs. Table 3 reports the results obtained with the Tiny-BioMoE model trained from scratch, retaining the same architecture but without pretraining. Table 4 shows the corresponding results using the pretrained version of Tiny-BioMoE, as described in Section 3.3, where the model acts as an embedding backbone. In all cases, the model was fine-tuned during training rather than kept frozen. For the BVP modality, training from scratch resulted in an accuracy of $45.24\%$ with the Angle representation, $39.30\%$ for Phase, and $42.58\%$ for PSD. The Recurrence plot reached $42.09\%$, while the highest performance was achieved with the Scalogram at $66.53\%$. The Waveform also performed reasonably well with $42.90\%$ accuracy. Using the pretrained model led to improvements across most representations. The most notable improvement was observed for Recurrence, which increased by over $5\%$, reaching $67.13\%$. In the case of EDA, overall performance was higher. With the scratch-trained model, Angle achieved $63.02\%$, and Scalogram reached $73.41\%$, again emerging as the top representation. Surprisingly, pretraining led to a slight decrease in performance: $61.36\%$ for Angle and $71.15\%$ for Scalogram. Notably, PSD achieved a precision of $80.60\%$, one of the highest observed. Respiration signals also showed strong results, with $45.61\%$ for Angle and $71.91\%$ for Scalogram. The pretrained model yielded minor fluctuations, with Scalogram improving to $73.53\%$ and Angle slightly dropping to $44.78\%$. For the SpO2 modality, pretraining resulted in consistent improvements across nearly all representations. Angle rose from $51.59\%$ to $55.80\%$, Phase from $48.12\%$ to $52.19\%$, and PSD from $47.86\%$ to $54.97\%$. The most substantial gains were observed for Recurrence and Waveform, which increased by $13.46\%$ and $15.62\%$, respectively. The only exception was Scalogram, which slightly decreased to $74.31\%$. Overall, the average accuracy across all representations and modalities increased from $52.13\%$ with the scratch-trained model to $53.96\%$ with the pretrained one. The most significant improvement occurred in the SpO2 modality, where the average accuracy rose from $55.59\%$ to $62.73\%$. Note that all the following experiments report results using the pretrained version of Tiny-BioMoE.
Table 3. Comparison of performance across different modalities and representations without pretraining (scratch).
Modality Representation Task–MC
Accuracy Precision F1
BVP Angle 45.24 49.42 46.72
BVP Phase 39.30 61.30 40.48
BVP PSD 42.58 58.24 45.03
BVP Recurrence 42.09 45.95 43.47
BVP Scalogram 66.53 67.33 66.41
BVP Waveform 42.90 43.71 41.86
EDA Angle 63.02 69.12 64.83
EDA Phase 53.94 58.26 53.21
EDA PSD 53.06 43.17 47.06
EDA Recurrence 59.37 63.27 61.11
EDA Scalogram 73.41 73.59 73.03
EDA Waveform 61.48 68.32 63.45
Resp Angle 45.61 51.85 45.57
Resp Phase 40.96 45.11 42.85
Resp PSD 43.42 57.49 46.08
Resp Recurrence 36.16 39.08 36.22
Resp Scalogram 71.91 71.89 71.72
Resp Waveform 36.62 40.20 32.90
SpO2 Angle 51.59 58.85 53.76
SpO2 Phase 48.12 72.53 51.26
SpO2 PSD 47.86 67.33 50.23
SpO2 Recurrence 55.18 59.81 57.06
SpO2 Scalogram 75.93 77.24 75.63
SpO2 Waveform 54.87 59.74 56.03
Table 4. Comparison of performance across different modalities and representations with the pretrained model.
Modality Representation Task–MC
Accuracy Precision F1
BVP Angle 46.06 49.71 47.40
BVP Phase 40.39 44.43 41.44
BVP PSD 43.04 48.00 44.42
BVP Recurrence 47.43 49.80 48.26
BVP Scalogram 67.13 65.42 62.19
BVP Waveform 43.67 51.28 44.00
EDA Angle 61.36 63.19 61.85
EDA Phase 51.29 58.22 53.87
EDA PSD 52.57 80.60 50.82
EDA Recurrence 55.48 57.05 56.22
EDA Scalogram 71.15 71.83 69.98
EDA Waveform 63.46 61.16 62.00
Resp Angle 44.78 45.80 45.19
Resp Phase 40.47 46.28 41.73
Resp PSD 45.29 51.40 46.18
Resp Recurrence 38.18 39.12 38.47
Resp Scalogram 73.53 73.93 73.40
Resp Waveform 33.33 26.78 29.70
SpO2 Angle 55.80 58.72 56.15
SpO2 Phase 52.19 56.37 53.65
SpO2 PSD 54.97 61.29 57.14
SpO2 Recurrence 68.64 70.45 69.18
SpO2 Scalogram 74.31 74.37 74.25
SpO2 Waveform 70.49 71.66 71.03
4.2. Fusion of Representations
The next series of experiments focuses on the fusion of representations within each modality. As previously discussed, each biosignal modality includes six distinct visual representations; however, not all contribute equally to performance. The goal is to evaluate which combinations of these representations can enhance the results. Fusion is performed by first extracting embeddings for each representation using the pretrained Tiny-BioMoE, followed by applying two fusion methods: element-wise addition and concatenation. Two main strategies are examined: (i) fusion of all six representations for each modality, and (ii) fusion of the two top-performing representations per modality. The results are presented in Table 5. For the BVP modality, fusion of all six representations using addition yielded an accuracy of $54.01\%$, while concatenation improved the performance to $58.87\%$, the highest recorded for this modality. In contrast, fusion of only the Scalogram and Recurrence representations resulted in lower accuracy: $53.08\%$ with addition and $49.18\%$ with concatenation. In all cases, fusion led to decreased performance compared to using only the Scalogram, the best standalone representation for BVP. Regarding the EDA modality, fusion of all representations decreased accuracy to $67.75\%$ with addition and $63.62\%$ with concatenation. However, when only Scalogram and Recurrence were fused, accuracy improved to $76.85\%$ and $77.88\%$ respectively—both outperforming the best individual representation for this modality. A similar trend was observed in the respiration signals. Fusion of all representations via concatenation resulted in $69.87\%$, which was lower than the $73.53\%$ achieved using only the Scalogram. Finally, for the SpO2 modality, fusion of all six representations did not improve performance. However, fusing Scalogram and Recurrence via concatenation led to an accuracy of $74.54\%$, showing a slight improvement over the best single-representation result.
Table 5. Comparison of performance across different fusion methods of representations.
Modality Representation Fusion Task–MC
Accuracy Precision F1
BVP Scal, Recur, Angle, Wave PSD, Phase add 54.01 61.22 56.62
BVP Scal, Recur, Angle, Wave PSD, Phase concat 58.87 63.40 60.18
BVP Scal, Recur add 53.08 54.20 53.06
BVP Scal, Recur concat 49.18 61.13 53.28
EDA Scal, Wave, Angle, Recur PSD, Phase add 67.75 68.01 64.63
EDA Scal, Wave, Angle, Recur PSD, Phase concat 63.62 61.59 61.64
EDA Scal, Recur add 76.85 73.70 75.06
EDA Scal, Recur concat 77.88 76.34 76.88
Resp Scal, PSD, Angle, Phase Recur, Wave add 66.18 70.30 65.27
Resp Scal, PSD, Angle, Phase Recur, Wave concat 69.87 67.91 68.81
Resp Scal, PSD add 67.36 71.34 64.90
Resp Scal, PSD concat 69.10 69.90 69.17
SpO2 Scal, Wave, Recur, Angle PSD, Phase add 71.96 74.59 72.84
SpO2 Scal, Wave, Recur, Angle PSD, Phase concat 73.84 74.11 73.79
SpO2 Scal, Recur add 72.92 72.93 72.90
SpO2 Scal, Recur concat 74.54 74.95 74.22
4.3. Fusion of Modalities
The final experiments investigate the fusion of modalities along with their corresponding representations. Initially, the best performing representation for each modality is selected and fused using both addition and concatenation methods, as in previous experiments. Additionally, a configuration using only the Scalogram representation from each modality is examined, since it generally yields the highest individual performance. The results are shown in Table 6. Specifically, the selected configuration includes the Scalogram for BVP, the Scalogram and Waveform for EDA, the Scalogram for respiration, and the Scalogram and Waveform for SpO2. This fusion setup resulted in high performance, achieving $81.02\%$ accuracy with addition and $82.41\%$ with concatenation—the highest scores reported in this study. Using only the Scalogram from each modality for fusion did not further improve performance, but still yielded strong results, with both fusion methods reaching an accuracy of $74.77\%$. Figure 5 visualizes the accuracy performance across individual modality representations, fused representations, and multimodal combinations reported in this study.
Table 6. Comparison of performance across different fusion methods of modalities.
Modality Representation Fusion Task–MC
Accuracy Precision F1
BVP Scalogram add 81.02 81.72 81.36
EDA Scal, Wave
Resp Scalogram
SpO2 Scal, Wave
BVP Scalogram concat 82.41 84.25 82.43
EDA Scal, Wave
Resp Scalogram
SpO2 Scal, Wave
BVP Scalogram add 74.77 75.47 74.84
EDA Scalogram
Resp Scalogram
SpO2 Scalogram
BVP Scalogram concat 74.77 75.75 75.23
EDA Scalogram
Resp Scalogram
SpO2 Scalogram
Figure 5. Performance across modality representations, combinations of representations, and combinations of modalities.
5. Comparison with Existing Methods
In this section, the proposed approach is compared with previous studies using the testing set of the AI4PAIN dataset. Some of these studies were conducted as part of the First Multimodal Sensing Grand Challenge. In contrast, others, including the present work, utilized data from the Second Multimodal Sensing Grand Challenge. The key distinction between the two challenges lies in the set of available modalities. Studies employing facial video or fNIRS have reported strong results, with accuracies of $49.00\%$ by (Prajod et al., 2024) and $55.00\%$ by (Nguyen et al., 2024), respectively. Combining these two modalities also yielded competitive results, although not substantially better than using each modality alone. For instance, (Vianto et al., 2025) reported $51.33\%$, and (Gkikas et al., 2025c) achieved $55.69\%$ using fused video and fNIRS data. Regarding the physiological modalities available in the Second Grand Challenge, some approaches achieved notably high performance, while others produced more limited results. In (Gkikas et al., 2025b), an accuracy of $55.17\%$ was reported using EDA in isolation, whereas (Gkikas et al., 2025a) achieved $42.17\%$ using only the respiration signal. The proposed method, which combines all available modalities—EDA, BVP, respiration, and SpO2—and utilizes specific visual representations for each, achieved an accuracy of $54.89\%$. This ranks among the highest performances reported on the dataset.
Table 7. Comparison of studies on the testing set of the AI4Pain dataset.
Study Modality ML Parameters Acc (%)
(Khan et al., 2025a) fNIRS ENS 53.66
(Nguyen et al., 2024) fNIRS Transformer 55.00
(Prajod et al., 2024) Video 2D CNN 49.00
(Gkikas and Tsiknakis, 2024b) Video, fNIRS Transformer 32.92 46.67
(Vianto et al., 2025) Video, fNIRS CNN-Transformer 86.73 51.33
(Gkikas et al., 2025c) Video, fNIRS Transformer 29.45 55.69
(Gkikas et al., 2025b) EDA Transformer 19.60 55.17
(Gkikas et al., 2025a) Respiration Transformer 3.62 42.24
Our EDA, BVP, Resp, SpO2 MoE 7.34 54.89
[…]
6. Conclusion
This study presents our contribution to the Second Multimodal Sensing Grand Challenge for Next-Generation Pain Assessment (AI4PAIN), where all available modalities—EDA, BVP, respiration, and SpO2 were utilized. We introduced Tiny-BioMoE, a lightweight embedding model for biosignal analysis. With only $7.3$ million parameters, a computational cost of $3.04$ GFLOPs, and pretrained on $4.4$ million biosignal-based representations, it serves as an efficient and versatile solution for a wide range of physiology-related tasks. Evaluated on the AI4PAIN dataset for pain recognition, Tiny-BioMoE demonstrated strong performance across modalities. A variety of visual representations were also explored. Results showed that the pretrained version of Tiny-BioMoE consistently outperformed its non-pretrained counterpart, especially in cases where the individual modality alone had limited discriminative power. This is particularly important, as it enables the use of modalities that might otherwise be disregarded. A comprehensive set of experiments revealed a clear trend in performance, progressing from isolated representations to intra-modality fusion and ultimately to multimodal combinations. The highest accuracy was achieved through a multimodal approach, combining all four modalities and selected visual representations. We argue that small, efficient pretrained models, such as Tiny-BioMoE, are highly valuable and that future efforts should prioritize their development and open distribution to help democratize access to advanced physiological modeling, regardless of hardware constraints.
Safe and Responsible Innovation Statement
This work relied on the AI4PAIN dataset (Fernandez Rojas et al., 2023b, 2024a), made available by the challenge organizers, to assess automatic pain recognition methods. All participants confirmed the absence of neurological or psychiatric conditions, unstable health issues, chronic pain, or regular medication use during the session. Before the experiment, participants were thoroughly informed of the procedures, and written consent was obtained. The original study’s human-subject protocol received ethical clearance from the University of Canberra’s Human Ethics Committee (approval number: 11837). The proposed method was developed for continuous pain monitoring, aiming to enhance pain assessment protocols and improve patient care. However, as validation and testing were performed on controlled laboratory data, its deployment in real-world clinical settings requires further investigation and comprehensive evaluation.
Acknowledgements
This work has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No $945175$ (Project: CARDIOCARE).